Math of finance

Bob is paying 1184.87 in monthly mortgage payments. She wants to refinance her existing mortgage loan of $100,000 at 14% interest for 30 years that she obtained 4 years ago. Her mortgage officer informed her that he could get her a rate of 10% but the finance cost would include paying a prepayment penalty on the existing loan, which is equal to 6 months interest on the balance of the loan, plus a closing cost of $2000. Would it be worth if for Bob to refinance? Justify your answer.
Answer #1

First we will calculate the Outstanding loan balance of Bob loan:

Outstanding Balance = P* [(1 + r) - (1+r) (1 + r) - 1

where, P = Loan Amount = $ 100,000

r = periodic Interest = 0.14/12 = 0.01166

n= no of periods = 30 years*12 months = 360

m = no of periods lapsed = 4 years *12 = 48

[(1 + 0.01166)300 – (1 +0.01166) 481 Outstanding Balance = 100000* (1 +0.01166) 300 - 1

Outstanding Balance = 100000* 65.084661 – 1.745007 65.084661 - 1

Outstanding Loan Balance = $ 98837.46

- Monthly payment = $1184.87

No of years left on loan = 30- 4 = 26 years

Total Interest on these 26 years= Total Payment in 26 years - current Loan balance

= ($1184.87*26 years*12 months)-$98837.46

= $ 270841.98

- Prepayment Penalty =6 months Interest on loan balance = $ 98837.46*14%*6/12

= $ 6918.62

Closing Cost = $ 2000

cost associated to refinance = $ 6918.62 + $ 2000 = $ 8918.62

Now, calculating the monthly payment of refinance loan:

EMI =P*r* (1+r) (1+r) -1

where,P = Loan amount = $ 98837.46

r = periodic Interest = 0.10/12 = 0.00833

n= no of periods = (30-4) years*12 months = 312

EMI = 98837.46 * 0.00833 * (1 +0.00833) 312 (1 + 0.00833)312 - 1

EMI = 98837.46 * 0.00833 * 13.319465 13.319465 - 1

EMI = $ 890.50

Total interest over period of loan = ($ 890.50*312) - $ 98837.46

= $ 178,998.54

Total cost of new loan = $ 178998.54 + $8918.62

= $ 187917.16

As total cost of new loan is less than old loan, Bob should refinance.

If you need any clarification, you can ask in comments.

If you like my answer, then please up-vote as it will be motivating